# Mass spectrum of the $\Omega\bar{\Omega}$ states

### v1

@@ I Introduction
In the present study, we perform a systematic investigation of the $\Omega\bar{\Omega}$ baryonium states with quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, $0^{++}$, and $1^{++}$ within the framework of QCD sum rules (QCDSR). QCDSR offers a powerful nonperturbative tool that bridges the fundamental dynamics of QCD with hadronic phenomenology. Originally developed by Shifman, Vainshtein, and Zakharov Shifman , this methodology has been extensively and successfully applied to the study of both conventional and exotic hadrons Albuquerque:2013ija ; Wang:2013vex ; Govaerts:1984hc ; Reinders:1984sr ; P.Col ; Narison:1989aq ; Tang:2021zti ; Qiao:2014vva ; Qiao:2015iea ; Tang:2019nwv ; Wan:2020oxt ; Wan:2022xkx ; Zhang:2022obn ; Wan:2024dmi ; Tang:2024zvf ; Li:2024ctd ; Zhao:2023imq ; Yin:2021cbb ; Yang:2020wkh ; Wan:2024cpc ; Wan:2024pet ; Wan:2024ykm ; Zhang:2024jvv ; Tang:2024kmh ; Tang:2016pcf ; Tang:2015twt ; Qiao:2013dda ; Qiao:2013raa ; Wan:2020fsk ; Wan:2025xhf ; Chen:2014vha ; Azizi:2019xla ; Wang:2017sto ; Chung:1981wm . In Sec.II, we provide a concise overview of the QCD sum rule framework and introduce the key formulas employed in our calculations. In Sec.III, the masses of the $\Omega\bar{\Omega}$ baryonium states with quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, $0^{++}$, and $1^{++}$ are evaluated within the QCDSR approach. Section IV is devoted to an analysis of the possible decay channels of these $\Omega\bar{\Omega}$ baryonium states. Finally, a brief summary and concluding remarks are presented.
@@ II Formalism
$S^{q}_{jk}(x)\!\!$ $=$ $\!\!\frac{i\delta_{jk}x\!\!\!/}{2\pi^{2}x^{4}}-\frac{\delta_{jk}m_{q}}{4\pi^{2}x^{2}}-\frac{it^{a}_{jk}G^{a}_{\alpha\beta}}{32\pi^{2}x^{2}}(\sigma^{\alpha\beta}x\!\!\!/+x\!\!\!/\sigma^{\alpha\beta})-\frac{\delta_{jk}}{12}\langle\bar{q}q\rangle+\frac{i\delta_{jk}x\!\!\!/}{48}m_{q}\langle\bar{q}q\rangle-\frac{\delta_{jk}x^{2}}{192}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle$ (11)
$+$ $\frac{i\delta_{jk}x^{2}x\!\!\!/}{1152}m_{q}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle-\frac{t^{a}_{jk}\sigma_{\alpha\beta}}{192}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle+\frac{it^{a}_{jk}}{768}(\sigma_{\alpha\beta}x\!\!\!/+x\!\!\!/\sigma_{\alpha\beta})m_{q}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle\;.$
@@ III Numerical analysis
The determination of an appropriate continuum threshold $s_{0}$ is carried out following the methodology outlined in Refs.Qiao:2013dda ; Tang:2016pcf ; Qiao:2013raa . The guiding principle is to identify the value of $s_{0}$ that produces an optimal stability window in the mass curve of the baryonium state, within which the extracted mass exhibits minimal sensitivity to variations in the Borel parameter $M_{B}^{2}$. In practice, $s_{0}$ is varied by $0.1$ GeV to determine its lower and upper bounds, and the resulting range is used to estimate the associated uncertainty in $s_{0}$ Wan:2020oxt ; Wan:2020fsk .
@@
The uncertainties in the present results primarily originate from the limited precision of the quark masses, vacuum condensates, and the continuum threshold parameter $\sqrt{s_{0}}$. For clarity and ease of reference, the corresponding values of the Borel parameters, continuum thresholds, and the resulting mass predictions are summarized in Table 1.
@@ Acknowledgments
This work was supported in part by Specific Fund of Fundamental Scientific Research Operating Expenses for Undergraduate Universities in Liaoning Province under Grants No. LJ212410165019 and the National Natural Science Foundation of China under Grants No. 12147214.

### v2

@@ I Introduction
In the present study, we perform a systematic investigation of the $\Omega\bar{\Omega}$ baryonium states with quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, $0^{++}$, and $1^{++}$ within the framework of QCD sum rules (QCDSR). QCDSR offers a powerful nonperturbative tool that bridges the fundamental dynamics of QCD with hadronic phenomenology. Originally developed by Shifman, Vainshtein, and Zakharov Shifman , this methodology has been extensively and successfully applied to the study of both conventional and exotic hadrons Albuquerque:2013ija ; Wang:2013vex ; Govaerts:1984hc ; Reinders:1984sr ; P.Col ; Narison:1989aq ; Tang:2021zti ; Qiao:2014vva ; Qiao:2015iea ; Tang:2019nwv ; Wan:2020oxt ; Wan:2022xkx ; Zhang:2022obn ; Wan:2024dmi ; Tang:2024zvf ; Li:2024ctd ; Zhao:2023imq ; Yin:2021cbb ; Yang:2020wkh ; Wan:2024cpc ; Wan:2024pet ; Wan:2024ykm ; Zhang:2024jvv ; Tang:2024kmh ; Tang:2016pcf ; Tang:2015twt ; Qiao:2013dda ; Qiao:2013raa ; Wan:2020fsk ; Wan:2025xhf ; Chen:2014vha ; Azizi:2019xla ; Wang:2017sto ; Chung:1981wm ; Wan:2025bdr . In Sec.II, we provide a concise overview of the QCD sum rule framework and introduce the key formulas employed in our calculations. In Sec.III, the masses of the $\Omega\bar{\Omega}$ baryonium states with quantum numbers $J^{PC}=0^{-+}$, $1^{--}$, $0^{++}$, and $1^{++}$ are evaluated within the QCDSR approach. Section IV is devoted to an analysis of the possible decay channels of these $\Omega\bar{\Omega}$ baryonium states. Finally, a brief summary and concluding remarks are presented.
@@ II Formalism
$S^{q}_{jk}(x)\!\!$ $=$ $\!\!\frac{i\delta_{jk}x\!\!\!/\penalty 50}{2\pi^{2}x^{4}}-\frac{\delta_{jk}m_{q}}{4\pi^{2}x^{2}}-\frac{it^{a}_{jk}G^{a}_{\alpha\beta}}{32\pi^{2}x^{2}}(\sigma^{\alpha\beta}x\!\!\!/\penalty 50+x\!\!\!/\penalty 50\sigma^{\alpha\beta})-\frac{\delta_{jk}}{12}\langle\bar{q}q\rangle+\frac{i\delta_{jk}x\!\!\!/\penalty 50}{48}m_{q}\langle\bar{q}q\rangle-\frac{\delta_{jk}x^{2}}{192}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle$ (11)
$+$ $\frac{i\delta_{jk}x^{2}x\!\!\!/\penalty 50}{1152}m_{q}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle-\frac{t^{a}_{jk}\sigma_{\alpha\beta}}{192}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle+\frac{it^{a}_{jk}}{768}(\sigma_{\alpha\beta}x\!\!\!/\penalty 50+x\!\!\!/\penalty 50\sigma_{\alpha\beta})m_{q}\langle g_{s}\bar{q}\sigma\cdot Gq\rangle\;.$
@@ III Numerical analysis
The determination of an appropriate continuum threshold $s_{0}$ is carried out following the methodology outlined in Refs.Qiao:2013dda ; Tang:2016pcf ; Qiao:2013raa . The guiding principle is to identify the value of $s_{0}$ that produces an optimal stability window in the mass curve of the baryonium state, within which the extracted mass exhibits minimal sensitivity to variations in the Borel parameter $M_{B}^{2}$. In practice, $s_{0}$ is varied by $0.1$ GeV to determine its lower and upper bounds, and the resulting range is used to estimate the associated uncertainty in $s_{0}$ Wan:2020oxt ; Wan:2020fsk . In Appendix B, we discuss in detail the significance of the continuum threshold parameter $\sqrt{s_{0}}$ within the QCD sum rule framework and its impact on the extracted mass values.
@@
For clarity and ease of reference, the corresponding values of the Borel parameters, continuum thresholds, and the resulting mass predictions are summarized in Table 1.
The uncertainties in the present results primarily originate from the limited precision of the quark masses, vacuum condensates, and the continuum threshold parameter $\sqrt{s_{0}}$. Each parameter is varied independently within its accepted range, while keeping the others fixed at their central values. The resulting variation in the extracted mass is taken as the individual uncertainty associated with that parameter:
$\Delta m_{i}=|m(p_{i}^{max})-m(p_{i}^{min})|.$ (23)
Here, $p_{i}$ denotes the parameters. The total uncertainty is estimated by summing individual contributions in quadrature:
$\Delta m_{totle}=\sqrt{\Sigma_{i}\Delta m_{i}^{2}}.$ (24)
The numerical contributions from the continuum threshold, quark mass, and condensate values are summarized in Table 2. From the table, one can see that the dominant sources of uncertainty originate from the condensate values and the continuum threshold.
@@
Table 2: Individual $\Delta m_{i}$ (GeV) contributions to the uncertainties of the extracted masses.
$J^{PC}$ from $s_{0}$ from $m_{s}$ from condensates
[…]
$1^{--}$ $\pm 0.05$ $\pm 0.02$ $\pm 0.06$
[…]
$1^{++}$ $\pm 0.08$ $\pm 0.03$ $\pm 0.07$
@@ V Summary
It should be noted that the strong interaction in the $\Omega\bar{\Omega}$ system is expected to exhibit important similarities to that in the $\Omega\Omega$ system, in view of their closely related quark content and the underlying QCD dynamics. For the $\Omega\Omega$ ground state, lattice QCD calculations predict a binding energy of about 32 MeV in Ref. Dhindsa:2025gae , about 7 MeV in Ref. Yamada:2014jra , and about 1.6 MeV in Ref. Gongyo:2017fjb . These values show a slight discrepancy with our result, in which the binding energy of the $0^{-+}$ $\Omega\bar{\Omega}$ state is estimated to be approximately $(124.9\pm 70.0)$ MeV. It is worth emphasizing, however, that the $\Omega\bar{\Omega}$ system corresponds to a baryonium configuration, where quark–antiquark annihilation effects may play a significant role, in contrast to the $\Omega\Omega$ dibaryon system. Indeed, within the framework of the chiral quark model and including the annihilation effect, the binding energy was found to be about 37–130 MeV in Ref. Zhang:2006dy , which is consistent with our result.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 12575106 and 12147214, and Specific Fund of Fundamental Scientific Research Operating Expenses for Undergraduate Universities in Liaoning Province under Grants No. LJ212410165019.
Appendix A The spectral densities
The spectral densities, including contributions from the OPE, are presented below.
[…]
$+s^{4}\left(70560m_{s}^{6}+4704m_{s}^{4}s+112m_{s}^{2}s^{2}+s^{3}\right)$
[…]
Here, $O3_{s}$, $O4$, and $O5_{s}$ are shorthand for $\langle\bar{s}s\rangle$, $\langle G^{2}\rangle$, and $\langle\bar{s}Gs\rangle$, respectively.
Appendix B Role of the Continuum Threshold Parameter $\sqrt{s_{0}}$ in QCD Sum Rule Analysis
To examine the sensitivity of the extracted mass to the continuum threshold parameter, we vary $\sqrt{s_{0}}$ within a reasonable range around its central value. As shown in Fig. 5, although the mass exhibits a slight dependence on $\sqrt{s_{0}}$, a stable platform can still be observed in the chosen Borel window. This behavior indicates that our results are not overly sensitive to the choice of $\sqrt{s_{0}}$, and the corresponding uncertainty has been properly taken into account in the final error estimation.
Figure 5: In the figure, the curves from bottom to top correspond to $\sqrt{s_{0}}$ values of 3.2, 3.3, 3.4, 3.5, 3.6, 3.7, and 3.8 GeV, respectively..
For the $0^{-+}$ state, as shown in Fig. 5, the curve is the flattest when $\sqrt{s_{0}}=3.5$ GeV, indicating the weakest dependence of the mass on the Borel parameter. When $\sqrt{s_{0}}=3.3$ GeV, the curve shows a noticeable downward trend, while for $\sqrt{s_{0}}=3.7$ GeV, the curve exhibits a clear upward trend. Therefore, we choose $\sqrt{s_{0}}=3.5$ GeV.
A similar analysis has been performed for the other states, with the corresponding figures shown in Fig. 6, Fig. 7, and Fig. 8, which correspond to the $1^{--}$, $0^{++}$, and $1^{++}$ states, respectively.
Figure 6: In the figure, the curves from bottom to top correspond to $\sqrt{s_{0}}$ values of 3.3, 3.4, 3.5, 3.6, 3.7, 3.8, and 3.9 GeV, respectively..
Figure 7: In the figure, the curves from bottom to top correspond to $\sqrt{s_{0}}$ values of 3.5, 3.6, 3.7, 3.8, 3.9, 4.0, and 4.1 GeV, respectively..
Figure 8: In the figure, the curves from bottom to top correspond to $\sqrt{s_{0}}$ values of 3.6, 3.7, 3.8, 3.9, 4.0, 4.1, and 4.2 GeV, respectively..